# Tensor network influence functionals in the continuous-time limit: connections to quantum embedding, bath discretization, and higher-order time propagation

### v1

@@ IV.1 Effective bath orbitals versus bath discretization
Figure 7: The difference in double occupancy $\langle n_{\uparrow}n_{\downarrow}\rangle$ between two different time-steps, $\Delta t$ and $2\Delta t$ using discrete-time IF-MPS with the second- and 4th-order Trotter decomposition (labeled as Trotter2 and Trotter4 with red triangles and green squares, respectively) and continuous-time IF-MPS with the 4th order Runge Kutta method and the second-order Trotter decomposition (labeled as RK4 and cont-Trotter2 with blue circles and cyan diamonds, respectively) at propagated times, $\Gamma t=2,3,$ and $5$, with the number of effective bath orbitals, $N_{\text{eff}}=8$. The results are fitted by the function, $\log|\langle n_{\uparrow}n_{\downarrow}\rangle(2\Delta t)-\langle n_{\uparrow}n_{\downarrow}\rangle(\Delta t)|=m\log\Delta t+b$, and the fits are displayed with dashed lines.
@@
As described in the analysis in Secs. II.4, the IF-MPS time propagation can be viewed as propagating a wavefunction in the Liouville space of the impurity and a set of effective bath orbitals. We now compare this time-dependent discretization of the (Liouville) bath with a more standard time-independent discretization.
@@
In Fig. 6, we show the time-dependence of the double occupancy of the impurity, $\langle n_{\uparrow}n_{\downarrow}\rangle$ in the quench dynamics. We show (second order Trotter discrete-time) IF-MPS dynamics generated using effective bath orbital numbers $N_{\text{eff}}=4,6,8,10$ with time-step $\Gamma\Delta t=0.01$ (Fig. 6a). We also show results from a time-independent bath discretization scheme based on a Lanczos iteration with thermofield transformation [66, 67, 5]. The thermofield transformation transforms the initial thermal state with $N_{b}$ bath orbitals to a Fermi sea with $N_{b}$ filled orbitals and $N_{b}$ empty orbitals and the Lanczos tridiagonalization algorithm over the filled and empty orbital space gives a truncated bath basis. We denote the number of Lanczos vectors in the truncated thermofield bath as $N_{b}^{\text{th}}$. We show dynamics using truncated thermofield bath discretizations with $N_{b}^{\text{th}}=4,6,8,10$ in Fig. 6b. The thermofield bath dynamics was propagated using the quantum chemistry full configuration interaction method and 4th-order Runge-Kutta with time-step $\Delta t=0.005$.
@@
As expected, the time-independent bath discretization yields substantial finite-size errors due to the limited number of Lanczos vectors supported by the bath, while the effective bath orbitals encode much more faithful dynamics. Indeed, at $N_{\text{eff}}=8$, the dynamics is fully converged to the eye (the corresponding error from the reference DMRG dynamics is shown in the inset), while the $N^{\text{th}}_{b}=10$ dynamics fails for $\Gamma t>2$. This illustrates the compactness of the IF-MPS time-dependent bath description, which resembles the behavior seen in real-time quantum embedding studies [37] even though the bath dynamics in the current formulation is non-unitary.
@@ IV.2 Converging to the continuous-time limit
We now analyze the time-step error incurred by varying the time-step, $\Gamma\Delta t\in\{0.02/2^{n}|n=0,1,2,\cdots,6\}$, at three different propagated times, $\Gamma t=2$, $3$, and $5$. The number of effective bath orbitals is fixed as $N_{\text{eff}}=8$, which is well converged to the reference data as shown in the previous section. We examine the convergence by fitting to the function, $f(\Delta t)=A\Delta t^{m}+C$. The exponent $m$ is extracted from the difference, $f(2\Delta t)-f(\Delta t)=A(2^{m}-1)\Delta t^{m}\propto\Delta t^{m}$, without the knowledge of the constant term, $C$.
In Fig. 7, the differences in $\langle n_{\uparrow}n_{\downarrow}\rangle$, $|\langle n_{\uparrow}n_{\downarrow}\rangle(2\Delta t)-\langle n_{\uparrow}n_{\downarrow}\rangle(\Delta t)|$, are illustrated as a function of different time-steps, $\Delta t$. We compare four different boundary IF-MPS schemes to analyze the time-step errors - 1) discrete-time boundary IF-MPS with the second-order Trotter decomposition (Trotter2, Eq. II.1), 2) discrete-time boundary IF-MPS with the 4th-order Trotter decomposition (Trotter4, Eq. IV), 3) continuous-time IF-MPS propagated by the 4th-order Runge-Kutta (RK4) method, and 4) continuous-time IF-MPS with the second-order Trotter decomposition (cont-Trotter2), which will be defined further below. The exponent $m$ is extracted by fitting the differences at the four smallest time-steps, and the fitted function is plotted as dashed lines.
For all three different propagated times, the exponent for the RK4 method has 4th-order time-step errors, which agrees with the theoretical scaling. However, the exponent extracted from both the second-order and 4th-order Trotter decomposition in the discrete-time formulation shows that the error is first-order in time-step, despite the theoretical error scaling of each Trotter decomposition without tensor network compression. We further note that the extrapolations to the continuous-time limit ($\Delta t=0$) from Trotter2 and RK4 agree with each other up to $10^{-6}$, whereas the extrapolated results from the Trotter4 data do not (Supplemental Material). This is because the continuous-time limit of Trotter2 (for finite bond dimension) is the same as that of RK4, but that of Trotter4 is not because it involves time evolution in both the forward and backward time directions.
The above discrepancies in the time-step error scaling of the Trotter schemes imply that there is an additional first-order time-step error associated with the discrete-time tensor network compression. We can show that the first-order time-step error arises from the projection of the reference (i.e., $N_{b}=40$) bath orbital space into the embedding space. For simplicity, we demonstrate this using a time-independent projection operator $\mathcal{P}$ without the gauge transformation and $\hat{L}_{S}=0$,
@@
Now consider decreasing the time-step $\Delta t$ to $\Delta t/2$. Then the time evolution operator for $\Delta t$ can be written as
@@
The two expressions agree to first-order in the time-step, but not to second-order in the time-step, due to the difference between $\mathcal{P}\hat{L}^{2}\mathcal{P}$ and $(\mathcal{P}\hat{L}\mathcal{P})^{2}$. Thus, after summing over $t/\Delta t$ time-steps, the total propagation has first-order time-step error, $\mathcal{O}(\Delta t)$. The IF-MPS compression involves time-dependent projection operators $\mathcal{P}(t)$ together with gauge transformations $\hat{G}(t)$. However, by assuming the smoothness of $\mathcal{P}(t)$ and $\hat{G}(t)$ in time, it is easy to show that the discrete-time IF-MPS also has this first-order time-step error.
@@
where the integration is over the time-step interval $\Delta t$ and $\mathcal{T}$ refers to a fermionic time-ordering operator. This can also be interpreted as a generalized time-dependent Trotter decomposition [68] of the embedding Liouville operator. The time-ordered operator can be expressed using the following differential equation,
@@
and the desired operator is $\hat{U}(\Delta t)$. Since $\hat{L}^{\text{emb}}_{SB}$ is a noninteracting operator, solving this differential equation of motion for $\hat{U}(x)$ is efficient (noting that the time evolution operator at $x=0$ is given by the identity operator, $\hat{U}(0)=\hat{I}$). We solve the differential equation using 4th-order Runge Kutta with time-step $2\Delta t_{\text{fine}}$, which allows us to include all the time-information of $\hat{L}^{\text{emb}}_{SB}$ obtained from the bath 1-RDM propagation with a $\Delta t_{\text{fine}}$ time-step.
In Fig. 7, time-step errors from the second-order cont-Trotter decomposition of the continuous-time IF-MPS clearly show that this gives second-order time-step errors. Indeed, the time-step propagation error with this technique is better even than that of 4th-order Runge-Kutta except for very small time-steps.
@@ V Conclusions
In this work, by analyzing the tensor network influence functional (IF-MPS) for the Anderson impurity model, we derived the correct continuous-time limit via the boundary IF-MPS. We further established a correspondence to discretized bath dynamics and quantum embedding. These formal results clarify the connection between the IF-MPS and other long-standing numerical techniques. They also provide the foundation to develop improved numerical implementations, for example, through the higher-order propagators identified in this work.
The numerical results we obtained on the quench dynamics of the Anderson model demonstrate the advantages of the current formulation. For example, compared to discrete-time IF-MPS, by using the equations of motion we derive, we can obtain high-order convergence of the time-step error. This is in contrast to increasing the order of the Trotterization in the standard discrete time approach, which does not in fact improve the time-step convergence, due to the errors associated with compression.
@@
The connections between IF-MPS and standard state-vector propagation further open up the application of a wide variety of wavefunction-based simulation tools within the boundary IF-MPS framework. We plan to explore the potential of these developments in a variety of physical applications in future work.

### v2

@@ IV.1 Effective bath orbitals versus bath discretization
Figure 7: The difference in double occupancy $\langle n_{\uparrow}n_{\downarrow}\rangle$ between two different time steps, $\Delta t$ and $2\Delta t$ using discrete-time IF-MPS with the second- and 4th-order Trotter decomposition (labeled as Trotter2 and Trotter4 with red triangles and green squares, respectively) and continuous-time IF-MPS with the 4th order Runge Kutta method and the second-order Trotter decomposition (labeled as RK4 and cont-Trotter2 with blue circles and cyan diamonds, respectively) at propagated times, $\gamma t=2,3,$ and $5$, the inverse temperature, $\gamma\beta=2$, with the number of effective bath orbitals, $N_{\text{eff}}=8$. The results are fitted by the function, $\log|\langle n_{\uparrow}n_{\downarrow}\rangle|_{2\Delta t}-\langle n_{\uparrow}n_{\downarrow}\rangle|_{\Delta t}|=m\log\Delta t+b$, and the fits are displayed with dashed lines.
@@
As described in the analysis in Sec. II.4, the IF-MPS time propagation can be viewed as propagating a wavefunction in the Liouville space of the impurity and a set of effective bath orbitals. We now compare this time-dependent discretization of the (Liouville) bath with a more standard time-independent discretization.
@@
In Fig. 6, we show the time-dependence of the double occupancy of the impurity $\langle n_{\uparrow}n_{\downarrow}\rangle$ in the quench dynamics. We show (second-order Trotter discrete-time) IF-MPS dynamics generated using effective bath orbital numbers $N_{\text{eff}}=4,6,8,10$ with time step $\gamma\Delta t=0.01$ with two different temperatures, $\gamma\beta=2$ (Fig. 6a) and $\gamma\beta=50$ (Fig. 6c). We also show results from a time-independent bath discretization scheme based on a Lanczos iteration with thermofield transformation [72, 73, 9]. The thermofield transformation transforms the initial thermal state with $N_{b}$ bath orbitals to a Fermi sea with $N_{b}$ filled orbitals and $N_{b}$ empty orbitals and the Lanczos tridiagonalization algorithm over the filled and empty orbital space gives a truncated bath basis. We denote the number of Lanczos vectors in the truncated thermofield bath as $N_{b}^{\text{th}}$. We show dynamics using truncated thermofield bath discretizations with $N_{b}^{\text{th}}=4,6,8,10$ in Figs. 6b and 6d. The thermofield bath dynamics was propagated using the quantum chemistry full configuration interaction method and 4th-order Runge-Kutta with time step $\Delta t=0.005$.
@@
As expected, the time-independent bath discretization yields substantial finite-size errors due to the limited number of Lanczos vectors supported by the bath, while the effective bath orbitals encode much more faithful dynamics. In the intermediate-temperature regime ($\gamma\beta=2$), the dynamics is fully converged to the eye at $N_{\text{eff}}=8$, while the $N^{\text{th}}_{b}=10$ dynamics fails for $\gamma t>2$. In the low-temperature regime ($\gamma\beta=50$), the IF-MPS dynamics with $N_{\text{eff}}=10$ faithfully reproduces the reference result within an accuracy of $10^{-3}$. This illustrates the compactness of the IF-MPS time-dependent bath description, which resembles the behavior seen in real-time quantum embedding studies [44] even though the bath dynamics in the current formulation is nonunitary.
@@ IV.2 Converging to the continuous-time limit
We now analyze the time step error incurred by varying the time step, $\gamma\Delta t\in\{0.02/2^{n}|n=0,1,2,\cdots,6\}$, at three different propagated times, $\gamma t=2$, $3$, and $5$. For this analysis, the inverse temperature is fixed as $\gamma\beta=2$ and the number of effective bath orbitals is fixed as $N_{\text{eff}}=8$, which is well converged to the reference data as shown in the previous section. We examine the convergence by fitting to the function $f(\Delta t)=A\Delta t^{m}+C$. The exponent $m$ is extracted from the difference, $f(2\Delta t)-f(\Delta t)=A(2^{m}-1)\Delta t^{m}\propto\Delta t^{m}$, without the knowledge of the constant term $C$.
In Fig. 7, the differences in $\langle n_{\uparrow}n_{\downarrow}\rangle$ between two different time steps, $\Delta t$ and $2\Delta t$, $|\langle n_{\uparrow}n_{\downarrow}\rangle|_{2\Delta t}-\langle n_{\uparrow}n_{\downarrow}\rangle|_{\Delta t}|$, where $\langle\bullet\rangle|_{\Delta t}$ means that the numerical simulation was performed with a timestep $\Delta t$, are illustrated as a function of different time steps $\Delta t$. We compare four different boundary IF-MPS schemes to analyze the time step errors: 1) discrete-time boundary IF-MPS with the second-order Trotter decomposition (Trotter2, Eq. II.1), 2) discrete-time boundary IF-MPS with the 4th-order Trotter decomposition (Trotter4, Eq. 55), 3) continuous-time IF-MPS propagated by the 4th-order Runge-Kutta (RK4) method, and 4) continuous-time IF-MPS with the second-order Trotter decomposition (cont-Trotter2), which will be defined further below. The exponent $m$ is extracted by fitting the differences at the four smallest time steps, and the fitted function is plotted as dashed lines.
For all three different propagated times, the exponent for the RK4 method has 4th-order time step errors, which agrees with the theoretical scaling. However, the exponent extracted from both the second-order and 4th-order Trotter decomposition in the discrete-time formulation shows that the error is first-order in time step, despite the theoretical error scaling of each Trotter decomposition without tensor network compression. We further note that the extrapolations to the continuous-time limit ($\Delta t=0$) from Trotter2 and RK4 agree with each other up to $10^{-6}$, whereas the extrapolated results from the Trotter4 data do not (see the SM). This is because the continuous-time limit of Trotter2 (for finite bond dimension) is the same as that of RK4, but that of Trotter4 is not because it involves time evolution in both the forward and backward time directions.
The above discrepancies in the time step error scaling of the Trotter schemes imply that there is an additional first-order time step error associated with the discrete-time tensor network compression. We can show that the first-order time step error arises from the projection of the reference (i.e., $N_{b}=40$) bath orbital space into the embedding space. For simplicity, we demonstrate this using a time-independent projection operator $\mathcal{P}$ without the gauge transformation and $\hat{L}_{S}=0$,
@@
Now consider decreasing the time step $\Delta t$ to $\Delta t/2$. Then, the time-evolution operator for $\Delta t$ can be written as
@@
The two expressions agree to first-order in the time step but not to second-order in the time step due to the difference between $\mathcal{P}\hat{L}^{2}\mathcal{P}$ and $(\mathcal{P}\hat{L}\mathcal{P})^{2}$. Thus, after summing over $t/\Delta t$ time steps, the total propagation has first-order time step error $\mathcal{O}(\Delta t)$. The IF-MPS compression involves time-dependent projection operators $\mathcal{P}(t)$ together with gauge transformations $\hat{G}(t)$. However, by assuming the smoothness of $\mathcal{P}(t)$ and $\hat{G}(t)$ in time, it is easy to show that the discrete-time IF-MPS also has this first-order time step error.
@@
where the integration is over the time step interval $\Delta t$ and $\mathcal{T}$ refers to a fermionic time-ordering operator. This can also be interpreted as a generalized time-dependent Trotter decomposition [74] of the embedding Liouville operator. The time-ordered operator can be expressed using the following differential equation,
@@
and the desired operator is $\hat{U}(\Delta t)$. Since $\hat{L}^{\text{emb}}_{SB}$ is a noninteracting operator, solving this differential equation of motion for $\hat{U}(x)$ is efficient (noting that the time-evolution operator at $x=0$ is given by the identity operator $\hat{U}(0)=\hat{I}$). We solve the differential equation using 4th-order Runge Kutta with time step $2\Delta t_{\text{fine}}$, which allows us to include all the time information of $\hat{L}^{\text{emb}}_{SB}$ obtained from the bath 1-RDM propagation with a $\Delta t_{\text{fine}}$ time step.
In Fig. 7, time step errors from the second-order cont-Trotter decomposition of the continuous-time IF-MPS clearly show that this gives second-order time step errors. Indeed, the time step propagation error with this technique is better even than that of 4th-order Runge-Kutta except for very small time steps.
@@ V Conclusions
In this paper, by analyzing the tensor network influence functional (IF-MPS) for the Anderson impurity model, we derived the correct continuous-time limit via the boundary IF-MPS. We further established a correspondence to discretized bath dynamics and quantum embedding. These formal results clarify the connection between the IF-MPS and other long-standing numerical techniques. They also provide the foundation to develop improved numerical implementations, for example, through the higher-order propagators identified in this paper.
The numerical results we obtained on the quench dynamics of the Anderson model demonstrate the advantages of the current formulation. For example, compared to discrete-time IF-MPS, by using the equations of motion we derive, we can obtain high-order convergence of the time step error. This is in contrast to increasing the order of the Trotterization in the standard discrete time approach, which does not in fact improve the time step convergence, due to the errors associated with compression.
@@
The connections between IF-MPS and standard state vector propagation further open up the application of a wide variety of wavefunction-based simulation tools within the boundary IF-MPS framework. We plan to explore the potential of these developments in a variety of physical applications in future work including its applications to the nonequilibrium Kondo problems or nonequilibrium dynamical mean-field theory.